OpenStudy (anonymous):

How many gallons of 20 % of salt solution must be mixed with 6 gallonsof 30% salt solution to make 22% salt solution?

OpenStudy (anonymous):

.2x + .3(6) = .22(x + 6) .2x + 1.8 = .22x + 1.32 .48 = .02x x = 24 gallons .2(24) + 1.8 = .22(24 + 6) 4.8 + 1.8 = .22(30) 6.6 = 6.6